A sum of money is shared among three brothers. The eldest receives 7/13 of it and the next receives 2/3 of the remainder. If the youngest brother received $6 , find the sum of money shared.
Answer by Theo(13342)

let x equal the original amount of money.

the eldest receive 7/13 of it, which is 7/13 * x.

what remains is 6/13 * x.

the next eldest receives 2/3 of that.

what remains is 1/3 of that, which is 1/3 * 6/13 * x = 6/39 * x.

the youngest receives 6 dollars, which is presumably the rest of the money.

therefore, 6/39 * x = 6

solve for x to get:

x = 39 * 6 / 6 = 39.

the original sum of money is 39.

the eldest receives 7/13 of 39 = 21 dollars.

what is left is 18 dollars.

the next receives 2/3 of 18 dollars = 12 dollars.

what is left is 6 dollars.

that's what the youngest gets.

the solution looks good.

the solution is:

the sum of money shared is 39 dollars.